Find an equation of the line that passes through the two given points. Write the equation in slope-intercept form, if possible.
passes through (5, 1) and (-5, 0)

Answer by TimothyLamb(4379)

slope = m = dy/dx
m = (1 - 0)/(5 - -5)
m = 1/10
---
y - 1 = m(x - 5)
y = mx - 5m + 1
y = (1/10)x - 5(1/10) + 1
y = (1/10)x - (1/2) + (2/2)
---
answer:
y = (1/10)x + 1/2
y = 0.1x + 0.5
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php